OpenStudy (anonymous):

Evaluate the expression (5x^2)^3/(10y)^4 for x = 10 and y = 5.

OpenStudy (anonymous):

(5x^2)^3/(10y)^4 = (5*100)^3/(50)^4 = 500^3/50^4 = 50^3*10^3/50^4 = 10^3/50 = 1000/50 = 20.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome. Medal for courtesy please? ;)

OpenStudy (anonymous):

im new to the site. whats this whole medal thing?

OpenStudy (anonymous):

You click "Best Response" for appreciation.